OpenStudy (anonymous):

Evaluate -100^(-3/2)

OpenStudy (anonymous):

-100?

OpenStudy (anonymous):

no the answer is-1/1000 but I don't know how to get to there!!

OpenStudy (anonymous):

\[ -100^{-3/2} = -\frac{1}{\sqrt{100^3}} =-\frac{1}{\sqrt{10^6}} = -\frac{1}{10^3} = -\frac{1}{1000} \]